A generator is designed to produce a maximum emf of 120 V while rotating with an angular speed of 3600 rpm. Each coil of the generator has an area of 0.033 m2. If the magnetic field used in the generator has a magnitude of 0.050 T, how many turns of wire are needed?
Answer #1

Solution) Maximum emf , E = 120 V

Angular speed , w = 3600 rpm

rpm = revolutions per minute

1 revolution = 2(pi) radians

1 minute = 60 seconds

1 rpm = (2(pi))/(60) rad/s

w = 3600×(2(pi))/(60)

w = 376.99 = 377 rad/s approximately

Area A = 0.033 m^2

Magnetic field B = 0.050 T

Number of turns , N = ?

We have , Maximum emf E = BANw

N = (E)/(BAw)

N = (120)/(0.050×0.033×377)

N = 192.91 = 193 approximately

N = 193 turns